OpenStudy (anonymous):

Find 2 integers who difference is 24 and whose product is as small as possible.

jigglypuff314 (jigglypuff314):

oh this one's easier than the other ones xD :P so if x and y represent the two integers then x - y = 24 and Min = xy

jigglypuff314 (jigglypuff314):

from x - y = 24 you can get that y = x - 24 so M = x(x-24) = x^2 - 24x then take the derivative of M

OpenStudy (anonymous):

one sec jiggles

OpenStudy (anonymous):

2x-24?

jigglypuff314 (jigglypuff314):

then to minimize it 2x - 24 = 0 solve for x

OpenStudy (anonymous):

is it just 12?

jigglypuff314 (jigglypuff314):

x = 12 yup then plug into y = x - 24

OpenStudy (anonymous):

-12?

jigglypuff314 (jigglypuff314):

@sourwing am I doing this right?

jigglypuff314 (jigglypuff314):

yeah, that's what I got :)

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

yeah

jigglypuff314 (jigglypuff314):

ok, thanks :)